I need to write an equation in standard form of the line that passes through the given point and has the given slope.
(4,3),m=7

Answer by mananth(16946)

(4,3),m=7
..
y= mx +b
plug value of x,y & m
..
3=7*4+b
3=28+b
-28
b=-25
..
y= mx +b
m=7, b= -25
y = 7x-25 is the equation.
m.ananth@hotmail.ca